# Towards Clean-Label Backdoor Attacks in the Physical World

### v1

@@ 3.3 Data collection
Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 5 and Appendix C.4). The images are captured in various indoor/outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. Details of our dataset can be found in the Appendix C.
Table 1: Attack success rate (%) of different poisoning algorithms on 6 random source-target poison pairs using sunglasses as the physical trigger. The attacker model and the victim model are both ResNet50. The results are averaged over 10 runs, with standard deviation in the bracket.
Poison pair Poison ratio $\alpha$ Naive BLC FM GM $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ $\epsilon=8/225$ $\epsilon=16/225$ 1 – 7 5% 50.7 (30.7) 13.4 (10.3) 21.4 (16.2) 77.5 (13.4) 92.3 (0.0) 99.4 (1.0) 100.0 (0.0) 10% 57.3 (12.0) 25.4 (16.5) 18.5 (13.1) 93.4 (1.8) 95.4 (2.0) 99.4 (0.5) 100.0 (0.0) 3 – 6 5% 20.8 (27.2) 31.4 (9.6) 68.6 (4.4) 98.6 (1.4) 99.5 (0.8) 98.6 (1.4) 99.5 (0.8) 10% 42.5 (14.5) 38.6 (13.5) 66.2 (5.5) 99.0 (0.8) 100.0 (0.0) 99.0 (11.3) 100.0 (0.0) 5 – 1 5% 0.9 (0.8) 0.0 (0.0) 0.6 (1.1) 0.0 (0.0) 5.9 (3.6) 26.5 (14.0) 88.0 (4.0) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 13.9 (9.3) 33.3 (6.0) 96.6 (1.9) 6 – 8 5% 1.8 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 32.5 (5.2) 31.1 (16.0) 98.1 (1.1) 10% 1.9 (1.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 37.2 (4.2) 43.5 (20.4) 95.6 (2.0) 7 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.7 (3.9) 46.6 (12.2) 98.5 (0.9) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.3 (0.4) 18.8 (4.6) 38.6 (15.5) 97.5 (1.1) 9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.5 (2.2) 9.7 (6.5) 4.8 (2.0) 85.9 (1.8) 10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 5.8 (5.9) 31.0 (8.1) 25.3 (11.3) 94.9 (2.3)
@@ 4 Experiments
We evaluate CLPBA using four Convolutional Neural Networks (CNNs), but only the results for ResNet50 [26] are presented in the main paper. The models are pre-trained on the VGGFace 2 dataset [27] and then fine-tuned on our dataset. We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. Each attack has a unique set of parameters, so we follow as closely as possible to the experimental settings of the original works. We measure the performance of an attack, or attack success rate (%) (ASR) success rate, by the percentage of trigger samples in the source class that are misclassified as the target class. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% and is therefore not reported in most experiments. Detailed experimental settings can be found in Appendix D. Additional experiments on other models and different settings are provided in Appendix E.
@@ 4.1 Comparison of poisoning algorithms
The results from Table 1 show that GM is the most effective poisoning algorithm. BLC and Naive methods fail with small poison ratios. Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values. Variance of experiment results between poisoning attempts could be due to different weight initialization of the victim model.
  • Why does GM work? By aligning the gradients of poisoned samples with the gradient of adversarial loss, we observe that the perturbations of GM appear to reconstruct deep representations of $\widetilde{\mathcal{D}}_{s}$. These perturbations replicate the facial contours of the source class and the appearance of the sunglasses trigger. This enables the model to extract key features of $\widetilde{\mathcal{D}}_{s}$ via the perturbations.
  • Why does FM fail? FM was initially proposed for transfer-learning models, assuming the weights of previous layers remain unchanged. In the end-to-end scenario, when the victim model is trained on the poisoned dataset, its shallower convolution kernels will learn the perturbations as features of the target class, making poisoned instances stay in the target-class distribution in deeper layers.
  • Why does BLC fail? First, the use of a low poison ratio and a subtle digital trigger (opacity 32/255) fails to establish a strong association between the trigger and the target class so the presence of the trigger at inference time cannot dominate the salient features in the source class. Second, the digital trigger may lack the variability introduced by physical conditions such as brightness changes, color shifts, noise, and viewpoint distortions, making it an unsuitable replacement for the physical trigger.
Among the three poisoning algorithms, FM produces the finest and most naturally dispersed perturbations (Appendix F.3); in contrast, BLC’s perturbations are most noticeable with irregular color patches all over the images. Interestingly, GM’s perturbations appear to reconstruct the face shape of the source class and the appearance of the triggers (Figure 3). In a small number of FM’s poisoned samples, the perturbations exhibit similar patterns to GM (e.g., face shape and textures of sunglasses). We believe that such samples are critical samples for the attack success because they hold the trigger information and the robust features of the source class. GM is more effective than FM since it can craft a greater number of critical samples.
4.2 Impact of poison pair on attack performance
In Table  1, the attack effectiveness varies significantly across source-target class pairs. Poison pairs $1-7$ and $3-6$ show much higher ASR than others, even for Naive and BLC poisoning methods. This section examines the impact of source and target classes on attack performance.
From the source class. A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, which enables the trigger features to outweigh salient features and cause misclassification. In Table 2, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for the source class 3 than for the source class 4, and thus the attack on class 3 has a higher ASR. In Figure 8 of Appendix B, we use Grad-CAM [28] to visualize discriminative regions implicitly detected by a clean ResNet50 model (i.e., model trained solely on clean data). We can see that in trigger instances of source class 3, the model struggles to extract discriminative features and must rely on less prominent features for classification. If the model is backdoored, then the model will focus on the trigger for classification.
Poison pair Clean model loss Poisoned model ASR (%)
@@ 4.2 Impact of poison pair on attack performance
From the target class. A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features. For example, in Appendix B, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger. The existence of such support samples helps enforce the attacker’s trigger as a dominant classification feature of the target class once the model is poisoned. Filtering these support samples greatly reduces the victim model’s susceptibility to the backdoor trigger (see Table 6, Appendix B). This naturally implies the existence of natural physical backdoors in a dataset, especially if unique physical objects exist in target class data. A detailed analysis of the natural backdoors is given in Appendix B. We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
4.3 Trigger selection
The performance of a backdoor attack has been shown to depend on the size, shape, and location of a trigger [10, 29]. The larger and closer to the face the trigger is, the more effectively it can obscure a person’s facial features and become dominant features. For a physical backdoor attack targeted at a facial recognition system, however, it is not practical to use large triggers like face masks and sunglasses, as people may be requested to take off such items when they go through facial recognition processes. Thus, it is important to study the performance of CLPBA using more stealthy triggers.
Table 3: Performance of backdoor attacks using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$. Triggers are arranged from left to right based on their stealthiness.
@@ 4.3 Trigger selection
We evaluate GM and FM attacks on poison pairs $7-3$ and $9-6$ using triggers of different sizes and levels of stealthiness (see Appendix C.2 for details into our trigger selection). We observe that the effectiveness of a trigger depends on the poison pair. Surprisingly, in poison pair $9-5$, the face mask, a large trigger, is not as effective as the "fake beard" trigger, a small and stealthy trigger. Factors that impact trigger effectiveness may include (1) whether the trigger aligns with the natural features of the target class and (2) whether the trigger hides important features of the source class.
A natural question for the attacker is "how to select an effective and stealthy trigger for a given poison pair?" The attacker can brute-force search for an effective trigger by iteratively crafting poisoned samples and evaluating the ASR for each of the triggers he can use. However, this could be too computationally and time-expensive if the dataset size or the number of triggers is large. Here, we propose a simple heuristic for choosing an effective trigger. Given a set of $k$ triggers $T=\{p_{1},p_{2},\ldots p_{k}\}$ corresponding to $\{\widetilde{D}_{s}^{(1)},\widetilde{D}_{s}^{(2)},\ldots\widetilde{D}_{s}^{(k)}\}$, in which $\widetilde{D}_{s}^{(i)}$ is the set of attacker’s source class samples with trigger $p_{i}$. With a model $F_{\theta}$ that has been trained only on clean data, the attacker can select a trigger for a poison pair $(s,t)$ with the following formula:
$p=\operatorname*{arg\,min}_{i}\frac{1}{|\widetilde{D}_{i}|}\sum_{x\in\widetilde{D}_{s}^{(i)}}2Softmax(F(x))_{t}-Softmax(F(x))_{s}$ (8)
where $Softmax(F(x))_{t}$ and $Softmax(F(x))_{s}$ represent the prediction probabilities of $x$ being in classes $t$ and $s$. Intuitively, Equation 8 is designed to identify a trigger that aligns with the natural features of the target class while obscuring key features of the source class. The constant in the first term is to bias our selection toward the first factor; $Softmax$ function is used to normalize the score for each sample. See Appendix H for an evaluation of our trigger selection heuristics.
4.4 Enhancing imperceptibility of attacks
In the pixel space. Higher values of $\epsilon$ result in higher ASR at the expense of visual stealthiness as the perturbations could leave behind visible artifacts on high-resolution facial images (Figure 3(a)). In the current $\ell_{\infty}$ regularisation, the projection of perturbations on the norm ball is only executed after every optimization step, making backdoor and visual imperceptibility two separate tasks. As a result, pixel values of the perturbations are often pushed to the boundary of the norm ball, causing noise to appear in unnecessary regions. Furthermore, the perturbations are non-smooth and irregular because each pixel is optimized individually. To improve the perceptual quality of poisoned samples, we propose to replace the current $\ell_{\infty}$ regularisation with a Pixel Regularization (PR) containing soft pixel-wise $\ell_{\infty}$ loss and Total Variation Upwind (TVU) [30]:
$\mathcal{R}_{pix}(\delta)$ $=||\delta-M\odot\epsilon||_{1}+TVU(\delta;\beta=1)$ (9)
where $M$ is a pixel-wise embedding mask based on the original image: $M_{i,j,k}=1.0-x_{i,j,k}$, assuming the pixel value $x_{i,j,k}\in[0,1]$. In Equation 10, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. Both terms in Equation 9 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels and $M$ keeping the perturbation within the color range. As shown in Figure 3, the perturbations crafted with PR are smoother and less conspicuous than those crafted with the original $\ell_{\infty}$ regularization. Following previous works [31, 32], we also evaluate the imperceptibility of poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [33]. The average PSNR value of poisoned samples is reported in Table 4.4. Our regularization results in higher PSNR, corresponding to better perceptual quality. We note that an additional advantage of the soft $\ell_{\infty}$ regularization over the standard one is that it is differentiable, which enables better convergence for the poisoning objective, especially for lower values of $\epsilon$ (see Appendix E.3).
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
Figure 3: Comparision between standard $\ell_{\infty}$ regularization with our Pixel Regularization. Perturbations are scaled 10×10$\times$10 × for visibility. The perturbations in both regularizations appear to reconstruct the facial shape of the source class with sunglasses as the trigger features.
In the feature space. Backdoor poisoning attacks have been observed to leave discernible traces in the feature space of poisoned models [34, 35, 36, 37]. By projecting feature representations of target samples (containing both clean and poisoned samples) on the first two principle axes (Figure 4(a)), we can see a clear distinction between the poisoned samples and clean samples. The separation exists because the perturbations on poisoned samples contain unnatural trigger features of $\widetilde{D}_{s}$, from which the model can learn as features of the target class. To limit the separation of poisoned samples from clean samples in the feature space, we propose a Feature Regularization (FR) that minimizes the distance of samples in $P$ to the distribution of original target-class samples $\mathcal{D}_{t}$: $\mathcal{R}_{feat}(\delta)=||f(x+\delta)-f(u)||_{2}^{2}$ where $u$ the empirical mean is computed from dataset $D_{t}$. In Figure 4, we can see that FR constrains the separation between poisoned and benign samples of the target class. In Table 4.4, by evaluating four representative filtering defenses (Spectral Signature (SS) [34], Activation Clustering (AC) [36], SPECTRE [35], DeepKNN [38]), we can confirm the effectiveness of FR in concealing the poisoned inputs from machine inspection in the feature space  [^2] Details of these defenses are outlined in Appendix G..
(a) Without FR
(b) With FR
Figure 4: 2D-PCA view of target-class representation produced by the poisoned model. FR constrains the separation between poisoned samples and benign samples.
Table 4: Defenses on GM attack ($\alpha=0.1,\epsilon=16/226$, poison pair $7-0$). ASR (%) is reported. PSNR (dB) is the metric to measure visual quality.
GM
GM + PR
GM + FR + PR
No defense 88.0 93.2 89.5
AC 0.0 0.9 88.9
SS 0.0 1.9 53.2
SPECTRE 0.0 99.1 87.0
DeepKNN 75.9 98.1 88.0
PSNR (dB) $\uparrow$ 29.8 dB 36.9 dB 37.2 dB
Altogether, the final form of poisoning objective is
$\min_{\delta}\mathcal{A}+\lambda_{1}\mathcal{R}_{pix}+\lambda_{2}\mathcal{R}_{feat}$ (11)
where $\mathcal{A}$ is the main poisoning objective, and $\lambda_{1},\lambda_{2}$ are balancing terms for the regularizations. In our experiments, we set $\lambda_{1}=\frac{10}{n}$ ($n$ is the total number of pixels), and $\lambda_{1}=0.05$. We note that $R_{feat}$ can degrade attack performance by constraining poison optimization, so only a small weight should be assigned to this regularization to mitigate the impact of filtering defenses.

### v2

@@ 3.3 Data Collection
Our facial dataset contains 21,238 images of 10 volunteers (9 males, 1 female, all Asian) and 7 triggers. It consists of 3,344 clean images, 9,790 trigger images of 7 physical triggers, and 8,104 additional images to evaluate accidental backdoor activations (see Section 7). Our animal dataset is created by combining a public Kaggle dataset [1], containing 45 classes of mammals, with a new cat class containing 500 cat images. For this dataset, we collected 1,500 cat images with 3 physical triggers (tennis ball, phone, book) to study CLPBA. The images are captured in various indoor and outdoor settings, under varying weather conditions, and with diverse shooting angles and distances. Details of our dataset can be found in the Appendix B.
Table 1: ASR (%) of poisoning algorithms on 5 random source-target pairs, averaged over 10 runs with standard deviation in brackets.
Poison pair Poison ratio $\alpha$ Naive MP FM GM
$\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$ $\epsilon=8/225$ $\epsilon=16/255$
(a) Facial Recognition (sunglasses trigger)
1 – 7 5% 35.0 (18.9) 14.2 (2.6) 10.5 (3.6) 84.0 (3.9) 80.9 (7.4) 97.2 (0.5) 99.4 (0.5)
10% 81.5 (3.5) 13.7 (2.3) 13.4 (2.6) 82.3 (0.5) 86.9 (2.5) 98.0 (0.5) 100.0 (0.0)
2 – 0 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 85.0 (4.4)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 99.4 (1.0)
4 – 3 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.4 (1.2) 15.8 (16.4) 96.4 (1.2)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 12.8 (11.6) 48.9 (16.2) 98.6 (0.3)
6 – 8 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.7 (2.9) 18.2 (12.3) 7.2 (11.0) 97.8 (2.4)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 4.7 (8.1) 22.9 (8.3) 21.8 (19.2) 98.3 (0.8)
9 – 5 5% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.6 (0.5) 5.8 (3.2) 22.7 (10.9) 95.6 (1.8)
10% 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 3.4 (0.5) 30.4 (1.5) 51.4 (17.8) 97.8 (0.2)
(b) Animal Classification (tennis trigger)
cat – fox 5% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.5) 0.2 (0.3) 8.2 (4.4) 9.6 (4.1)
10% N/A 0.0 (0.0) 0.0 (0.0) 1.0 (0.8) 0.2 (0.4) 20.5 (8.9) 17.8 (7.7)
cat – kangaroo 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.9 (0.1) 8.4 (0.8)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.1 (0.1) 0.1 (0.1) 9.4 (2.1) 18.0 (2.0)
cat – mongoose 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 2.7 (0.3) 5.3 (0.4)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 6.7 (0.1) 10.3 (2.9)
cat – sea-lion 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 7.4 (0.6) 13.4 (2.9)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.3 (0.1) 0.0 (0.0) 22.1 (5.3) 23.6 (5.8)
cat – tapir 5% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 1.8 (0.8) 7.8 (5.0)
10% N/A 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 29.8 (8.6) 36.3 (7.6)
@@ 4 Experiments
For Facial recognition, we choose the victim model to be ResNet50 [18] pre-trained on the VGGFace2 dataset [4]. For Animal classification, the victim model is ResNet18 pre-trained on Imagenet [43]. We assume a gray-box scenario where the attacker knows the architecture of the victim model but does not have access to its training process. To enable a fair comparison across different poisoning algorithms, we standardize the poisoning hyperparameters, including the number of iterations, optimizer, batch size, learning rate, and other relevant parameters. The performance of attacks is measured by attack success rate (%) (ASR), the percentage of trigger source-class samples that are misclassified as the target class. Note that the prediction accuracy (ACC) on clean samples consistently exceeds 99% for facial recognition and 93% for animal classification across different attacks, so it is not reported in the paper. Detailed experimental configurations can be found in Appendix C. Appendix D provides additional experiments on other architectures, hyperparameters, and physical triggers.
@@ 4.1 Comparison of Poisoning Algorithms
The results from Table 1 show that GM is the most effective poisoning algorithm. Higher values of $\epsilon$ and $\alpha$ could increase ASR, but $\epsilon$ has a greater impact as it allows the poisoning objective to converge to smaller values and the poisoned samples to contain more adversarial features of the source-class trigger distribution. Additionally, CLPBA is more effective for facial recognition than animal classification, probably because the ”sunglasses” is a very dominant feature that hides important classification features (i.e., the eyes and the nose) for face recognition, while the ”tennis ball” trigger is a relatively minor feature so it cannot outweigh the natural features of the cat.
Dataset bias. Interestingly, we can observe that the poison pair 1-7 in Table 1a has an unusually high ASR compared to other pairs. Even with no trigger images injected, ASR reaches 26.21 % for the Naive method. We believe this is because certain samples in the target class contain similar features to the physical trigger, reinforcing the trigger as a dominant feature once trigger images are added. In this case, it may be the eyeglasses of the target-class that have a similar shape and texture to the sunglasses. While this correlation may be accidental due to our dataset collection and trained classifier, it still raises an important issue of dataset bias that a backdoor attacker can exploit to select an effective trigger aligning with the natural features of the target class. In Appendix A, we show that the removal of training samples with strong trigger features drastically reduces the ASR and thus prevents the dataset bias.
4.2 Analysis of CLPBA
Two unique attributes of CLPBA are that it is hidden-trigger and source-specific. The first attribute means the trigger is not present in the poisoned samples, while the second ensures only source-class samples can cause the model to misclassify with the trigger. CLPBA achieves these attributes by embedding features of the source-class trigger distribution, $\widetilde{\mathcal{D}}_{s}$, into the perturbations in the poisoned target samples. Specifically, the embedded features consist of adversarial features and trigger features. We define adversarial features as the non-robust features [25] that shift the representations of target instances toward the source-class distribution in the feature space, and trigger features as the high-level features of the physical trigger extracted by the neural network. The existence of adversarial features ensures the source-specific attribute, while reconstructed trigger features allow the model to indirectly learn the physical trigger as a feature of the target class.
Figure 3: Experiment to show the existence of adversarial features. We set a high $\epsilon$ of 32, $\alpha=2$, poison pair is $9-5$.
Existence of adversarial features. To justify the existence of adversarial features, we take inspiration from [25] to test whether the model trained on poisoned target samples could successfully predict source-class samples. We design the experiment as illustrated in Figure 3. First, we craft the GM-poisoned samples with the original dataset and the attacker model. Then, we filter out all source-class samples, before relabelling all GM-poisoned samples to the source class. We then train the victim model from scratch on this modified dataset before evaluating the recall (true positive rate) of predicting source-class samples that have been filtered previously. To avoid randomness, we set a prediction threshold of 85% after the softmax function. Notably, the model trained on this modified dataset has a prediction accuracy of 75% on source-class samples. This shows that the perturbations on poisoned target instances contain adversarial features representative of source-class distribution. To make sure that the trained model is not biased towards predicting out-of-distribution samples as the source class, we also evaluate the false positive rate of classifying 200 random face images from the CelebA dataset [35] as the source class, and the accuracy is 0%.
Existence of trigger features. By scaling the perturbations on GM-poisoned samples, we observe that the perturbations appear to reconstruct the shape of the physical trigger. Interestingly, setting a high value of $\epsilon$ even reconstructs the samples in $\widetilde{\mathcal{D}}_{s}$, as can be observed in Figure 5.
(a) Clean model
(b) Poisoned model
Figure 4: Feature space visualization.
Feature space analysis. In Figure 4, we project source-class samples (green), target-class samples (blue), trigger source samples (yellow), and poisoned target samples (red) in 2-dimension feature space of both clean and poisoned models where the first axis is $\mu_{s}-\mu_{t}$ and the second axis is $\theta_{s}-\theta_{t}$, the normal vector to the decision boundary between the source and target class. In both feature spaces, we can see that some poisoned target instances are close to the samples of $\widetilde{\mathcal{D}}_{s}$. For the clean model, $\widetilde{\mathcal{D}}_{s}$ overlaps with ${\mathcal{D}}_{s}$ in the feature space. For the poisoned model, however, $\widetilde{\mathcal{D}}_{s}$ is shifted towards $\widetilde{\mathcal{D}}_{t}$ in the feature space. From this observation, we can say that since poisoned target instances contain representative features of $\widetilde{\mathcal{D}}_{s}$, especially the trigger feature which is not a part of ${\mathcal{D}}_{s}$, training the model on these poisoned instances cause $\widetilde{\mathcal{D}}_{s}$ to move closer to the classification region of the target class in the feature space.
Figure 5: From left to right: tennis, fake-beard, and sunglasses triggers. First row: a sample in $\widetilde{D}_{s}$. Second row: GM perturbation.
4.3 Defenses
We evaluate the performance of CLPBA against representative filtering defenses (filter poisoned samples during training), and firewall defenses (filter trigger samples during inference). For the filtering defenses, we evaluate Elimination Rate (ER), the percentage of poisoned samples that are correctly filtered, and Sacrifice Rate (SR), the percentage of clean samples that are incorrectly filtered. For firewall defenses, we evaluate the True Positive Rate (TPR), the percentage of trigger source-class samples that are correctly filtered, and the False Positive Rate (FPR), the percentage of non-trigger samples that are incorrectly filtered. For details on these defenses, we refer readers to Appendix F.
Table 2: Defense evaluation on GM attack ($\alpha=0.1$, $\epsilon=16/226$, poison pair $7-0$). ASR is 99.7% without defenses.
Filtering Defenses
Metrics AC[3] SS[52] SPECTRE[17] SCAN[51]
ER (%) 62.16 97.30 72.97 0
SR (%) 15.77 19.77 19.50 0
ASR (%) 99.7 0.0 99.7 99.7
Firewall Defenses
Metrics STRIP[14] CD[20] IBD-PSC[19] Frequency[59]
TPR (%) 0.0 79.6 0.0 0.0
FPR (%) 6.7 65.3 16.7 0.9
For filtering defenses, we evaluate Spectral Signature (SS) [52], Activation Clustering (AC) [3], SPECTRE [17], and SCAN [51]. Table 2 shows that only SS is effective against CLPBA. Although we can see that AC and SPECTRE have high elimination rates (62.16% and 72.97%), the attack still achieves 99.7 % ASR. This shows that the GM poisoning objective (Equation 8) focuses on a small subset of effective poison instances to match the gradient of adversarial loss and distill the features from the trigger distribution. This is supported in Figure 4 where only a subset of poisoned samples overlap with the trigger distribution. Detecting such effective instances is generally difficult since they correlate strongly with each other and they transform the model’s decision boundary of the target class. Table 2 also shows that firewall defenses are not effective for CLPBAs because they either miss trigger samples (low TPR) or incorrectly filter out too many benign samples (high FPR). This behavior is expected since these defenses are built on the memorization property of backdoor attacks, which differs from the generalization property of CLPBA.
5 Improving the Effectiveness of CLPBA
Empirically, higher $\epsilon$ lead to a higher ASR since the poisoning objective can converge to a smaller value. However, the perceptual quality of poisoned samples is compromised since the poisoned images have high-frequency artifacts that are clearly visible (Figure 5). The drawbacks of the $\ell_{\infty}$ regularization on perturbations are that it is non-differentiable, and the projection on $\ell_{\infty}$-norm ball is only executed at the end of an optimization step, making backdoor and poison stealthiness two separate tasks. As a consequence, the perceptual quality of poisoned images is compromised because (1) the pixel values of perturbation are often pushed to the boundary of the norm ball, (2) the noise can appear in unnecessary regions of the image, and (3) the perturbations contain high-frequency artifacts. To improve the visual quality and convergence of the poisoning objective, we propose to replace the current $\ell_{\infty}$ regularization with a Pixel regularization containing a soft pixel-wise $\ell_{\infty}$ loss and discretized Total Variation Upwind (TVU) [6]:
$\mathcal{R}_{pix}(\delta)$ $=||\delta^{+}-M\odot\epsilon||_{1}+\alpha TVU(\delta;\beta=1)$ (10)
where $\delta^{+}$ takes the absolute value of every element in $\delta$, $M$ is a mask of pixel-wise embedding ratio: $M_{i,j,k}=\frac{x_{i,j,k}}{\max_{k}{x_{i,j,k}}}$, assuming $x_{i,j,k}\in[0,255]$. $M$ encourages the perturbation to align with the underlying image content. In Equation 11, $\mathcal{N}(p)$ are neighbouring pixels of pixel $p$. The weighting factor $\alpha$ is set to 5. Both terms in Equation 10 improve the smoothness of perturbations, with $TVU(\delta)$ minimizing the difference between adjacent pixels. The side-by-side comparison between standard $\ell_{\infty}$ regularization and our pixel regularization is given in Figure 6. Our pixel regularization leads to a better convergence of the poisoning objective (Equation 8) while simultaneously reducing the $\ell_{2}$ norm of the perturbations to a smaller value, indicating an improved perceptual quality. Consequently, the GM attack with pixel regularization achieves an ASR of $97.8(1.1)$, compared to $51.9(18.7)$ for the original GM attack under the same settings. We also evaluate the perceptual quality of the poisoned samples statistically using Peak Signal-to-Noise Ratio (PSNR) [24], following previous works [29, 31]. Our pixel regularization results in higher values of PSNR (36.9dB vs 29.8dB), corresponding to a better perceptual quality.
(a) Standard $\ell_{\infty}$ regularization
(b) Pixel Regularization
(c) Poisoning loss
(d) Perturbation norm
Figure 6: Comparison between standard $\ell_{\infty}$-regularization and our Pixel Regularization. Top row: visual comparison; Bottom row: convergence rate. Poisoned samples are crafted with GM attack ($\alpha=0.1$, $\epsilon=8/255$) on poison pair 77773333.
6 Analysis of Attack Factors
The results in Table 1 show that ASR varies between different poison pairs for a given trigger, while the results in Table 4 show that different triggers may have different performances for a given poison pair. This section examines the impact of poison pairs and the physical trigger on CLPBA.
Poison pair Clean model loss Backdoored model ASR (%)
@@ 6 Analysis of Attack Factors
Source class. A source class is more susceptible to physical attacks at inference if the trigger obscures the class’s discriminative classification regions, enabling the trigger features to outweigh salient features and cause misclassification. In Table 3, the classification loss of the clean ResNet50 model on inference-time trigger samples is much higher for source class 3 than for source class 4, resulting in a higher ASR for class 3. In Appendix A, we use Grad-CAM [47] to visualize discriminative regions detected by a clean and poisoned ResNet50 model (i.e., trained on clean and poisoned data). We observe that in trigger instances of source class 3, the clean model struggles to extract meaningful facial features and relies on less prominent features for classification. Conversely, the poisoned model focuses on the physical trigger for classification.
@@
Table 4: Performance of GM attack using various physical triggers. Poisoned samples are crafted with $\epsilon=8$ and $\alpha=0.1$.
Target class. A target class can be more susceptible to CLPBA due to the similarity between the class’s natural features and the trigger features. For example, in Appendix A, we show that the eyeglasses in some samples of the target class could represent weaker backdoor features for the attacker’s trigger. We note that there is also an unexplored relationship between the distribution of samples in the target class $\mathcal{D}_{t}$ and the distribution of trigger samples in the source class $\widetilde{\mathcal{D}}_{s}$ that allows for feature alignment of poisoned samples in high dimensional space.
@@
Physical trigger. We evaluate GM and FM attacks on poison pairs $7-3$ and $9-5$ using triggers of different sizes and levels of stealthiness (see Appendix B for details on our trigger selection). We observe that the effectiveness of a trigger depends on the poison pair. Surprisingly, in poison pair $9-5$, the ”face mask”, a large trigger, is not as effective as the ”fake beard” trigger, a small and stealthy trigger. From our previous analysis, there are two factors impacting trigger effectiveness: whether the trigger aligns with the natural features of the target class, and whether the trigger hides important features of the source class.